a) Hugh has opened a retirement fund account which pays 7 percent interest and requires
$5,000 annual deposits. Hugh will retire in 15 years and expects 10 years of retirement life.
What is the maximum annual retirement benefit Hugh can get during his retirement years?

b) To expand its operation, Sunbeam Ltd. has applied to the Lion Bank for a 3-year, $3,500,000
loan. Prepare a loan amortization table assuming 10 percent rate of interest.
1
Expert's answer
2016-05-12T08:37:02-0400
a) i = 7%, $5,000 annual deposits. If Hugh will retire in 15 years and expects 10 years of retirement life, then the maximum annual retirement benefit Hugh can get during his retirement years is FV per year = (P*((1+r)^n - 1)/r)/10 = (5000*(1.07^15 - 1)/0.07)/10 = $12,564.5 per year.

b) If Sunbeam Ltd. has applied to the Lion Bank for a 3-year, $3,500,000 loan, then a loan amortization table assuming 10 percent rate of interest will look like:
Year Beginning Balance Total Payment Principal Interest Ending Balance
1 3,500,000 1,552,833.33 1,166,666.66 386,166.67 2,333,333.33
2 2,333,333.33 1,552,833.33 1,166,666.66 386,166.67 1,166,666.66
3 1,166,666.66 1,552,833.33 1,166,666.66 386,166.67 0
Payment per year is P = (3,500,000*(1 + 0.1)^3)/3 = $1,552,833.33.
Interest payment per year is I = (3,500,000*(1 + 0.1)^3) - 3,500,000 = 4,658,500 - 3,500,000 = 386,166.67